# Improving Value-based Process Verifier via Structural Prior Injection

### v1

@@ Abstract
In the Large Language Model(LLM) reasoning scenario, people often estimate state value via Monte Carlo sampling. Though Monte Carlo estimation is an elegant method with less inductive bias, noise and errors are inevitably introduced due to the limited sampling. To handle the problem, we inject the structural prior into the value representation and transfer the scalar value into the expectation of a pre-defined categorical distribution, representing the noise and errors from a distribution perspective. Specifically, by treating the result of Monte Carlo sampling as a single sample from the prior ground-truth Binomial distribution, we quantify the sampling error as the mismatch between posterior estimated distribution and ground-truth distribution, which is thus optimized via distribution selection optimization. We test the performance of value-based process verifiers on Best-of-N task and Beam search task. Compared with the scalar value representation, we show that reasonable structural prior injection induced by different objective functions or optimization methods can improve the performance of value-based process verifiers for about 1$\sim$2 points at little-to-no cost. We also show that under different structural prior, the verifiers’ performances vary greatly despite having the same optimal solution, indicating the importance of reasonable structural prior injection.
@@ 1 Introduction
To address the first problem, people often use the mean-square error objective function and map the discrete estimated value into the contiguous space(Lu et al., 2024; Luo et al., 2024; Snell et al., 2024). The mean-square error objective function introduces the distance prior and shows that the distance between two different values can be measured by the square of the difference between the two values, which extends the discrete estimated values into the contiguous space. To address the second problem, people often use the temporal difference method to update the value of the current state using values of subsequent states iteratively, greatly reducing the sampling variance at the cost of a larger sampling quantity(Tsitsiklis & Roy, 1997; Gallici et al., 2024). However, in the LLM reasoning scenario, it’s expensive to extend the sampling quantity as each sample requires a completed rollout to the final state, which limits the application of Monte Carlo estimation as the annotation of the state value.
@@
To handle the problem, we propose the distance metric called Statistics-based Distance, which is used to measure the distance between two categorical distributions conditioned on the probability of sampling the corresponding category from the ground-truth target distribution. The distance metric can be used to judge the reasonableness of the prior structural information to be injected, which can therefore guide the optimization approach. We optimize the categorical distribution via different objective functions, including the mean-square error and the cross-entropy(HL) loss(Imani & White, 2018), showing that the reasonable structural prior can improve the performance of value-based process verifiers on different objective functions and different tasks for about 1$\sim$2 points at little-to-no cost. Through ablation studies, we also show that different structural prior can result in very different performances.
@@ 3.1 Regression: from Scalar to Categorical Distribution
where $f_{\theta}$ is the function mapping from state space $\{q\}\bigcup s_{t}\in\mathcal{S}$ to the contiguous space $y\in[0,1]$, $y$ is the estimated value from the current state to the outcome that ranges from 0 to 1. We name Eqn.11 as scalar regression.
@@
1:  Input: questions $Q$, answers $A$, states $S$, policy $\pi$, verifier $f_{\theta}$, Monte Carlo sampling size $k$
@@
6:     if Scalar regression, Eqn.11then
@@
8:     else if Expectation regression, Eqn.14then
@@ 4.1 Experimental Settings
Following previous research (Wang et al., 2024; Lightman et al., 2024), we evaluate value-based process verifiers based on their verification ability through best-of-$N$ sampling. Specifically, given a problem $p$, we sample $N$ candidate solutions from a generator. Then, the candidates are re-ranked according to the score generated by the given verifier. The candidate with the highest score will be treated as the best candidate and selected as the final solution. Finally, we compare the consistency of the final solution and ground-truth answer to determine whether the solution is correct. The statistical success rate will be reported. Following Snell et al. (2024), we also include the beam-search metric to evaluate the performance of value-based process verifiers at test time. Specifically, the number of beams $N$ and beam size $M$ will be set. Given a problem $p$, the verifier is required to score every step during the generation process. For each iteration, $N$ beams will generate $M$ next-step candidates individually, the verifier will select the top $N$ candidates from the $N*M$ beams as the beams for the next iteration, and then continue the iteration until finished. Like Best-of-N, we compare the consistency of the final solution and ground-truth answer. The statistical success rate will be reported.
@@
The generator in our experiments is based on LLemma-7b. We train it on MetaMATH for 3 epochs to get the generator. Based on the train split of MATH dataset, we construct the training dataset of the process verifier. To clarify, We use the generator to sample 15 solutions per problem. Following previous works(Lightman et al., 2024; Wang et al., 2024), we split each solution into steps by the pre-defined rule-based strategies (e.g. newline as the delimiters). For each step, we combine it with its previous steps to form an incomplete solution, then sample 8 rollouts to perform Monte Carlo estimation and annotate the state value. In general, we sample $15*8=120$ samples for each problem and the training dataset has 180k samples in total.
@@ Appendix A Related Work
Several works have replaced regression with classification to improve performance in different domains(Weiss & Indurkhya, 1995; Zhang et al., 2023). Notably, Imani & White (2018) proposed the HL-Gauss cross-entropy loss as the drop-in replacement of MSE loss, showing that by incorporating prior distribution information, the classification loss can be effective in regression tasks. The authors then performed experiments in several different distributions and found that the HL-Gaussian distribution can greatly improve the performance, while other distribution-learning approaches can result in even worse performance. Farebrother et al. (2024) further extended the classification optimization method to several more complex domains, showing the general benefits of performing cross-entropy on the categorical distribution for regression tasks. However, how to perform better categorical distribution modeling is still unclear. In the LLM reasoning scenario, we endow the categorical distribution with fixed bins the specific definition, explaining the difference between regression and classification from the perspective of prior structural information injection. We show that the regression loss can also benefit from suitable categorical distribution modeling, indicating that the categorical distribution modeling can achieve the general benefits compared with scalar regression methods.

### v2

@@ Abstract
[…]
@@ 1 Introduction
To address the first problem, people often use the mean-square error objective function and map the discrete estimated value into the contiguous space(Lu et al., 2024; Luo et al., 2024; Snell et al., 2024). The mean-square error objective function introduces the distance prior and shows that the distance between two different values can be measured by the square of the difference between the two values, which extends the discrete estimated values into the contiguous space. To address the second problem, people often use the temporal difference method to update the value of the current state using values of subsequent states iteratively, greatly reducing the sampling variance at the cost of a larger sampling quantity(Tsitsiklis and Roy, 1997; Gallici et al., 2024). However, in the LLM reasoning scenario, it’s expensive to extend the sampling quantity as each sample requires a completed rollout to the final state, which limits the application of Monte Carlo estimation as the annotation of the state value.
@@
[…]
@@ 3.1 Regression: from Scalar to Categorical Distribution
where $f_{\theta}$ is the function mapping from state space $\{q\}\bigcup s_{t}\in\mathcal{S}$ to the contiguous space $\in[0,1], yy$ is the estimated value from the current state to the outcome that ranges from 0 to 1. We name Eqn.11 as scalar regression.
@@
1:Input: questions $Q$, answers AA, states $S$, policy $\pi$, verifier $f_{\theta}$, Monte Carlo sampling size $k$
@@
6:  if Scalar regression, Eqn.11: then
@@
8:  else if Expectation regression, Eqn.14: then
@@ 4.1 Experimental Settings
Following previous research (Wang et al., 2024; Lightman et al., 2024), we evaluate value-based process verifiers based on their verification ability through best-of-$N$ sampling. Specifically, given a problem $p$, we sample $N$ candidate solutions from a generator. Then, the candidates are re-ranked according to the score generated by the given verifier. The candidate with the highest score will be treated as the best candidate and selected as the final solution. Finally, we compare the consistency of the final solution and ground-truth answer to determine whether the solution is correct. The statistical success rate will be reported. Following Snell et al. (2024), we also include the beam-search metric to evaluate the performance of value-based process verifiers at test time. Specifically, the number of beams $N$ and beam size $M$ will be set. Given a problem $p$, the verifier is required to score every step during the generation process. For each iteration, $N$ beams will generate $M$ next-step candidates individually, the verifier will select the top $N$ candidates from the NMN*M beams as the beams for the next iteration, and then continue the iteration until finished. Like Best-of-N, we compare the consistency of the final solution and ground-truth answer. The statistical success rate will be reported.
@@
The generator in our experiments is based on LLemma-7b. We train it on MetaMATH for 3 epochs to get the generator. Based on the train split of MATH dataset, we construct the training dataset of the process verifier. To clarify, We use the generator to sample 15 solutions per problem. Following previous works(Lightman et al., 2024; Wang et al., 2024), we split each solution into steps by the pre-defined rule-based strategies (e.g. newline as the delimiters). For each step, we combine it with its previous steps to form an incomplete solution, then sample 8 rollouts to perform Monte Carlo estimation and annotate the state value. In general, we sample 158=12015*8=120 samples for each problem and the training dataset has 180k samples in total.
@@ Appendix A Related Work
Several works have replaced regression with classification to improve performance in different domains(Weiss and Indurkhya, 1995; Zhang et al., 2023). Notably, Imani and White (2018) proposed the HL-Gauss cross-entropy loss as the drop-in replacement of MSE loss, showing that by incorporating prior distribution information, the classification loss can be effective in regression tasks. The authors then performed experiments in several different distributions and found that the HL-Gaussian distribution can greatly improve the performance, while other distribution-learning approaches can result in even worse performance. Farebrother et al. (2024) further extended the classification optimization method to several more complex domains, showing the general benefits of performing cross-entropy on the categorical distribution for regression tasks. However, how to perform better categorical distribution modeling is still unclear. In the LLM reasoning scenario, we endow the categorical distribution with fixed bins the specific definition, explaining the difference between regression and classification from the perspective of prior structural information injection. We show that the regression loss can also benefit from suitable categorical distribution modeling, indicating that the categorical distribution modeling can achieve the general benefits compared with scalar regression methods.